# Thunder-NUBench: A Benchmark for LLMs’ Sentence-Level Negation Understanding

### v3

@@ 4 Thunder-NUBench Dataset
Pre-processing.
We begin by extracting sentences from two primary corpora: (1) the Hover dataset (jiang2020hover), designed for multi-hop fact extraction and claim verification, and (2) the Wikipedia Summary dataset (scheepers2017compositionality), which contains concise summaries from English Wikipedia. We chose these datasets because their factual content and complex sentence structures are well-suited for developing a dataset aimed at understanding standard negation in complex, sufficiently lengthy sentences. Additionally, we automatically correct any grammatical errors and merge or split sentences as needed to create well-formed single-sentence units.
Generation.
We create two types of datasets from the pre-processed sentences: the sentence-negation pair dataset and the multiple choice dataset. In the sentence-negation pair dataset, each original sentence is paired with a manually crafted standard negation, as detailed in Section 3.3. In the multiple-choice dataset, each original sentence is presented with four options: a standard negation, a local negation, a contradiction, and a paraphrase. Each of them are described in Table 4. Together, these categories assess whether models truly understand semantic negation rather than relying on superficial cues.
Standard and local negation options are manually created rather than generated by LLMs. We have observed that LLMs often struggle to produce correct standard negations, frequently resulting in subclausal or local negations instead. They can also generate incorrect local negations, even when explicitly prompted to do otherwise. Since precise negation is essential to our benchmark, these options must be developed by humans to ensure the quality of the dataset. In contrast, contradiction and paraphrase options are initially created automatically using carefully designed prompts with the OpenAI API (OpenAI) and are then refined during the review process.
@@ 5.1 Evaluation Setup
We evaluate models under two common Multiple-Choice Question Answering (MCQA) evaluation settings: (1) a completion-based evaluation, where the model assigns probabilities to each candidate by appending it as a continuation of the prompt, and (2) an option-selection evaluation, where the model selects from labeled options (A/B/C/D). In the completion-based evaluation, we report performance using accuracy, while in the option-selection evaluation, we use exact match. To mitigate known issues, such as selection or position bias in the option-selection evaluation, we randomly shuffle the order of the options (using random seed 42). Details of the specific prompt templates and formatting can be found in Appendix O.
Models.
We evaluate two main groups of model sizes: those with 2-3 billion parameters and those with 7-8 billion parameters (team2024gemma; grattafiori2024llama; qwen2025qwen25technicalreport; jiang2023mistral7b). Each group includes both pretrained models and instruction-tuned models. To also examine larger models, we incorporate API models, which are assessed only in the option-selection setting  (achiam2023gpt; hurst2024gpt; TheC3). This is because recent APIs do not provide the log-probability outputs needed for completion-based evaluation.
  • 2-3B models:
    • Pretrained: gemma-2b, Llama-3.2-3B, Qwen2.5-3B
    • Instruction-tuned: gemma-1.1-2b-it, Llama-3.2-3B-Instruct, Qwen2.5-3B-Instruct
  • 7-8B models:
    • Pretrained: gemma-7b, Llama-3.1-8B, Mistral-7B-v0.3, Qwen2.5-7B
    • Instruction-tuned: gemma-1.1-7b-it, Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, Qwen2.5-7B-Instruct
  • API models: GPT-4o mini, GPT-4.1 mini, Claude Haiku 3.5
Zero-shot and few-shot evaluation.
For each model, we evaluate performance in both zero-shot and few-shot settings using the Language Model Evaluation Harness (eval-harness). In the few-shot scenario, we use examples from the demonstration set as in-context demonstrations. Results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and are reported for one, five, and ten examples from the demonstration set (1-shot, 5-shot, and 10-shot). We present the performance results on the test set for each model and prompt configuration.
Supervised fine-tuning.
We conduct Supervised Fine-Tuning (SFT) using the LLaMA-Factory framework (zheng2024llamafactory) on the Sentence-Negation Pair dataset from Thunder-NUBench. The dataset is formatted in the Alpaca instruction style (alpaca). To achieve parameter-efficient training, we apply Low-Rank Adaptation (LoRA) (hu2022lora) with a rank of 8, targeting all linear layers. The fine-tuning process is carried out for three epochs, using a batch size of 1, a gradient accumulation step of 8, cosine learning rate scheduling, and bfloat16 precision. After the SFT, we evaluate the model’s zero-shot performance to directly assess its ability to generalize from instruction tuning without being influenced by in-context examples. It is important to note that API models are not fine-tuned, as they are not compatible with the LLaMA-Factory framework.
Figure 3: Model performance on Thunder-NUBench. Circles (blue) represent the average performance of 2-3B models, squares (purple) indicate the average for 7-8B models, upward triangles (orange) signify the average of base models, and downward triangles (red) denote the average of instruction-tuned models. Stars (green) represent API models.
5.2 Model Performance on Thunder-NUBench
Figure 3 displays evaluation results from all models across three settings: the zero-shot baseline, the few-shot baseline, and the zero-shot after SFT using Thunder-NUBench. It summarizes the overall trends concerning model sizes, training configurations (including whether it is instruction-tuned), and evaluation settings. Complete results are reported in Appendix P.
In the completion-based evaluation, performance steadily improves from zero-shot to few-shot and further after SFT, with gaps between smaller and larger models narrowing after SFT. In the option-selection evaluation, performance also rises with more shots, but SFT yields smaller gains than few-shot prompting. Meanwhile, instruction-tuned models remain relatively stable across conditions. API models achieve the best overall performance, though their performance doesn’t consistently improve with more shots, except at 10-shot.
5.3 Analysis of Negation Understanding Performance
We analyze model errors to evaluate the ability of our models to differentiate standard negation from similar semantic variants. Each type of local negation in our dataset is explicitly labeled based on its sentence structure: relative clause, participle clause, compound sentence, and adverbial clause, as defined in Table 3.
Error
Rate
@@ 5.3 Analysis of Negation Understanding Performance
Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
@@
completion- based Llama- 3.1-8B Baseline zeroshot 0.562 70.62 21.33 8.05 25.64 30.84 64.29 43.87
5shot 0.393 83.87 14.31 1.81 20.83 22.40 47.62 45.81
SFT zeroshot 0.203 85.55 12.50 1.95 8.97 11.69 18.71 32.26
Llama- 3.1-8B- Instruct Baseline zeroshot 0.488 70.57 26.18 3.25 24.36 25.97 54.76 37.74
5shot 0.347 80.55 18.76 0.69 18.91 19.16 39.80 37.74
SFT zeroshot 0.244 86.04 12.34 1.62 12.18 12.66 28.23 33.87
option- selection Llama- 3.1-8B Baseline zeroshot 0.541 47.95 6.74 45.31 23.08 24.03 41.50 19.03
5shot 0.362 78.95 6.58 14.47 30.45 21.75 40.82 25.16
SFT zeroshot 0.444 58.79 27.34 13.87 20.83 20.13 37.41 20.65
Llama- 3.1-8B- Instruct Baseline zeroshot 0.256 76.53 15.11 8.36 16.67 18.83 19.05 23.23
5shot 0.269 74.04 21.83 4.13 20.83 16.88 25.17 19.35
SFT zeroshot 0.254 72.32 19.03 8.65 15.06 17.86 21.09 14.52
@@
Table 6: Error distribution and confusion analysis of pretrained and instruction-tuned Llama-3.1-8B models across various evaluation settings: zero-shot baseline, five-shot baseline, and SFT zero-shot.
@@
To identify which subtypes of local negation are most frequently confused with standard negation, we calculate the confusion rate. This is defined as the proportion of examples within each subtype where the model incorrectly selects the local negation option instead of the correct standard negation. For example, if 320 items are labeled as participle clause negation and the model incorrectly chooses the local negation option instead of the correct standard negation option in 32 of these cases, the confusion rate for participle clause negation would be 10%. Complete analysis results are provided in Appendix R.
@@
We focus on the results of Llama-3.1-8B and its instruction-tuned version, as shown in Table 6. In the completion-based setting, error rates generally decrease from zero-shot to 5-shot and continue to improve after SFT, with most errors concentrated in local negation options. Within local negation, compound sentences exhibit the highest confusion but also show the largest relative improvement after SFT.
In the option-selection setting, the base model demonstrates unusually high errors for paraphrase options in the zero-shot scenario (nearly 40%). Although these errors decrease after SFT, they are partly offset by increases in errors for local negation and contradiction options. The instruction-tuned model consistently achieves lower overall error states and confusion rates than the base version. Compared to the completion-based setting, the option-selection setting results in lower confusion rates for adverbial clause negation.
@@ 6 Conclusion
Although we built the dataset using two distinct sources (HoVer and Wikipedia summaries), both are derived from encyclopedic, formal domains, which may not fully represent the variety of sentence structures and informal language found in real-world use cases. Moreover, while all examples were systematically generated and reviewed, some bias may persist due to subjective decisions in the human curation process. We attempt to mitigate this through cross-checking by an independent group of authors, but some residual bias may remain.
Thunder-NUBench primarily focuses on standard (sentence-level) negation and its distinction from local negation, contradiction, and paraphrase. Other important negation phenomena, such as double negation and negative polarity items (NPIs), are not directly addressed in this benchmark. Our current focus is establishing a strong foundation for evaluating models’ understanding of standard negation. However, we aim to expand the evaluation to a broader range of negation phenomena in future work.
@@
This work does not involve the use of crowd-sourcing methods. Instead, all data included in the Thunder-NUBench benchmark was carefully reviewed by the authors to ensure quality, relevance, and adherence to ethical standards. The datasets and tools used for training and evaluation are publicly available and used in compliance with their respective licenses.
@@
The Thunder-NUBench dataset will be released under the Creative Commons Attribution Non-Commercial Share Alike 4.0, ensuring transparency, reproducibility, and accessibility for future research.
@@ Appendix A Typology of Contradiction
Contradictions in natural language can arise in diverse ways that go beyond simple negation. Following the typology of de2008finding, contradictions can be grouped into seven categories: antonymy, explicit negation, numeric mismatch, factive/modal inconsistencies, structural reversals, lexical incompatibilities, and conflicts based on world knowledge. These categories reflect the fact that contradiction covers a broader semantic scope than negation alone. Table 7 summarizes these types with definitions and examples.
@@
Table 7: Contradiction types from de2008finding. Contradiction covers a broader scope than negation.
@@ Appendix B Copular Verbs
Copular verbs, also known as linking verbs, are verbs that connect the subject of a sentence to a subject complement, which can be a noun, adjective, or other expression that describes or identifies the subject. Unlike action verbs, copular verbs do not express actions but rather states or conditions. The most common copular verb in English is "to be" in its various forms (am, is, are, was, were). Other examples include "seem," "appear," "become," "feel," "look," "sound," "taste," and "smell" when used to describe the subject’s state (hengeveld1986copular).
@@ Appendix C Negation of Implications
Negating implications presents challenges, as natural language intuitions often diverge from the rules of formal logic. Let’s say there is a conditional statement, "If I study hard, I will pass the bar exam." Formally, let $P$ denote "I study hard" and $Q$ denote "I will pass the bar exam." In classical logic, the conditional "if $P$, QQ" can be false only when $P$ is true and $Q$ is false. This implies that the negation of the conditional is $P\;\text{and}\;\text{Neg}(Q)$" ("I study hard and I won’t pass the bar exam,) while the conditional itself is equivalent to $\text{Neg}(P)\;\text{or}\;Q$" ("I don’t study hard or I will pass the exam) (nguyen2023negation).
Psychological studies confirm that people often accept both "if $P$, Neg(QQ)" ("If I study hard, I won’t pass the exam.") and "if Neg($P$), QQ" ("If I don’t study hard, I will pass the exam."). However, the former can be interpreted as $\text{Neg}(P)\;\text{or}\;\text{Neg}(Q)$", and the latter $P\;\text{or}\;Q$", both of which are not equivalent to the original statement’s negation, $P\;\text{and}\;\text{Neg}(Q)$". "if Neg($P$), Neg(QQ)" ("If I don’t study hard, I won’t pass the exam.") is not the correct negation as well, as it is equivalent to $P\;\text{or}\;\text{Neg}(Q)$(espino2012not).
@@ Appendix D Compound Sentences and Coordinating Conjunction
A compound sentence consists of two or more independent clauses joined by a coordinating conjunction. Each clause can stand alone, but they are combined to express related ideas (gleitman1965coordinating).
@@ Appendix F Double Negation
Double negation refers to the use of two forms of grammatical negation within a single sentence. In standard English, only one negative form should be present in a subject-predicate construction; the presence of two negatives is generally considered non-standard and often results in an unintended meaning. For example, while "He’s going nowhere" is correct, "He’s not going nowhere" is ungrammatical. Another example is "I won’t bake no cake," which combines verb negation ("won’t") with object negation ("no cake"), resulting in a grammatically incorrect construction (deprez2015double).
In English, certain double negation constructions convey affirmative meanings rather than intensifying negation, effectively paraphrasing the original positive statement (e.g., $\neg\neg p\approx p$(van1996litotes). This rhetorical device, known as litotes, often manifests in expressions such as "not bad," implying "good," or "not unhappy," implying "happy." Leveraging this phenomenon, we have generated paraphrase candidates for our dataset using such double negation patterns. For example,

### v4

@@ 4 Thunder-NUBench Dataset
Pre-processing.
We begin by extracting sentences from two primary corpora: (1) the Hover dataset (Jiang et al., 2020), designed for multi-hop fact extraction and claim verification, and (2) the Wikipedia Summary dataset (Scheepers, 2017), which contains concise summaries from English Wikipedia. We chose these datasets because their factual content and complex sentence structures are well-suited for developing a dataset aimed at understanding standard negation in complex, sufficiently lengthy sentences. Additionally, we automatically correct any grammatical errors and merge or split sentences as needed to create well-formed single-sentence units.
Generation.
We create two types of datasets from the pre-processed sentences: the sentence-negation pair dataset and the multiple choice dataset. In the sentence-negation pair dataset, each original sentence is paired with a manually crafted standard negation, as detailed in Section 3.3. In the multiple-choice dataset, each original sentence is presented with four options: a standard negation, a local negation, a contradiction, and a paraphrase. Each of them is described in Table 4. Together, these categories assess whether models truly understand semantic negation rather than relying on superficial cues.
Standard and local negation options are manually created rather than generated by LLMs. We have observed that LLMs often struggle to produce correct standard negations, frequently resulting in subclausal or local negations instead. They can also generate incorrect local negations, even when explicitly prompted to do otherwise. Since precise negation is essential to our benchmark, these options must be developed by humans to ensure the quality of the dataset. In contrast, contradiction and paraphrase options are initially created automatically using carefully designed prompts with the OpenAI API (OpenAI, 2025) and are then refined during the review process. Details of the GPT models and code used at each stage of the data generation process are provided in Appendix K.
@@ 5.1 Evaluation Setup
We evaluate models under two common Multiple-Choice Question Answering (MCQA) settings: (1) a completion-based evaluation, where the model assigns probabilities to each candidate by appending it as a continuation of the prompt, and (2) an option-selection evaluation, where the model selects from labeled options (A/B/C/D). To mitigate known issues such as selection or position bias in the option-selection setting, we randomly shuffle the order of the options (using random seed 42). In both settings, we use two instruction variants: a definition-based instruction (referred to as the definition instruction) and a step-by-step instruction (referred to as the detailed instruction). Details of the prompt templates and formatting are provided in Appendix M.
We report results for a diverse set of pretrained and instruction-tuned models, evaluated under zero-shot and few-shot settings with 1, 5, and 10 examples, as well as supervised fine-tuning (SFT). For SFT, models are trained on the sentence-negation pair dataset introduced in Thunder-NUBench. The full list of models, evaluation protocols, and fine-tuning details is provided in Appendix N.
5.2 Overall Performance and Key Findings
Effect of Model Size.
Performance tends to improve as model size increases, indicating that negation understanding benefits from larger models. The improvement is more pronounced in the option-selection evaluation than in the completion-based setting, with larger models showing higher accuracy gains (as seen in Figure 3). This suggests that option selection benefits more from larger models, since it requires comparing multiple candidate answers rather than generating a single continuation.
Figure 3: Zero-shot accuracy of instruction-tuned Qwen3 models under the detailed prompt setting.
Effect of Instruction Tuning.
Instruction tuning generally improves performance across most model families and evaluation settings. As shown in Figure 4, this improvement is observed in both completion-based and option-selection evaluations. We further observe that instruction tuning yields larger improvements under the detailed prompt setting than under the definition-based instruction. This suggests that instruction-tuned models follow step-by-step instructions more reliably, leading to larger gains in negation understanding.
Figure 4: Boxplots showing the distribution of zero-shot accuracy for pretrained and instruction-tuned models across evaluation settings, using both definition-based and detailed prompt instructions.
However, the effect of instruction tuning is not uniform across all models. In particular, some Qwen3 models exhibit comparable or higher performance in their pretrained versions, indicating that instruction tuning does not always guarantee gains for negation understanding.
Effect of Few-Shot Learning.
Overall, increasing the number of in-context demonstrations tends to improve performance across few-shot settings. This indicates that few-shot learning generally benefits sentence-level negation understanding. Figure 5 illustrates this trend for the Llama-3.1-8B-Instruct model under definition-based instruction. The initial introduction of demonstrations leads to a clear performance improvement, while the incremental benefits decrease as more examples are added.
This trend, however, is not uniform. For some API-based models under the detailed instruction, additional demonstrations do not always yield further gains. One possible reason is that the detailed instructions already provide strong guidance for API-based large models, and adding demonstrations may introduce noise or encourage unnecessary pattern following.
Figure 5: Mean accuracy of Llama-3.1-8B-Instruct under the definition instruction across different numbers of in-context demonstrations.
@@ 5.2 Overall Performance and Key Findings
Evaluation Formats Instruction Formats Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
@@
completion- based definition baseline zero-shot 0.536 73.82 23.96 2.22 25.00 31.82 58.50 48.71
5-shot 0.367 85.31 14.25 0.43 18.91 22.40 40.14 48.06
after SFT zero-shot 0.229 86.51 12.46 1.04 10.90 14.29 23.13 33.55
detailed baseline zero-shot 0.468 71.86 26.44 1.69 23.08 29.87 47.62 38.71
5-shot 0.347 84.70 14.84 0.46 19.55 23.05 37.76 41.29
after SFT zero-shot 0.211 87.97 10.90 1.13 8.97 13.31 21.09 33.23
option- selection definition baseline zero-shot 0.462 70.67 18.52 10.81 27.56 24.68 53.40 30.00
5-shot 0.332 81.62 11.22 7.16 25.64 18.83 40.48 27.42
after SFT zero-shot 0.356 71.94 15.37 12.69 25.00 19.81 44.56 17.10
detailed baseline zero-shot 0.486 71.45 14.85 13.70 33.01 25.97 56.12 29.03
5-shot 0.353 77.53 11.24 11.24 27.56 19.81 39.12 26.77
after SFT zero-shot 0.270 65.69 21.70 12.61 14.74 14.61 29.25 15.16
@@
Table 6: Error distribution and confusion analysis of Llama-3.1-8B-Instruct model across various evaluation settings.
Effect of Supervised Fine-Tuning.
Supervised fine-tuning (SFT) generally leads to improved zero-shot performance across most models, indicating that task-specific training further enhances negation understanding. The performance gains from SFT are more pronounced in the completion-based evaluation than in the option-selection setting, suggesting that the effects of fine-tuning may interact with the evaluation format. We observe that SFT does not degrade models’ general language abilities, as confirmed by evaluations on general benchmarks (Appendix P).
Detailed results for all models and training settings are provided in Appendix O. In addition, we conducted a human evaluation, showing that although humans perform well on average, distinguishing standard negation from closely related alternatives is not uniformly easy across participants. Detailed results are provided in Appendix R.
5.3 Error Analysis
We analyze model errors to evaluate the ability of our models to differentiate standard negation from similar semantic variants. Each type of local negation in our dataset is explicitly labeled based on its sentence structure, as defined in Table 3.
@@ 5.3 Error Analysis
We measure the confusion rate, defined as the proportion of examples within each subtype where the model incorrectly selects the local negation option instead of the correct standard negation. For example, if 320 items are labeled as participle clause negation and the model incorrectly chooses the local negation option instead of the correct standard negation option in 32 of these cases, the confusion rate for participle clause negation would be 10%. Complete analysis results are provided in Appendix Q.
@@
We focus on the results of Llama-3.1-8B-Instruct model as shown in Table 6. In the completion-based setting, error rates generally decrease from zero-shot to 5-shot and continue to improve after SFT, with most errors concentrated in local negation options. Within local negation, compound sentences exhibit the highest confusion but also show the largest relative improvement after SFT.
In the option-selection setting, errors are likewise dominated by local negation, while the model shows a higher tendency to confuse paraphrase options compared to the completion-based setting. Comparatively, the option-selection setting results in lower confusion rates for adverbial clause negation. In addition, performance improvements in the option-selection setting do not consistently follow the expected progression across training configurations. In some cases, SFT yields higher error rates than the few-shot baselines, and this pattern is observed in some other models as well.
@@ 6 Conclusion
Although we built the dataset using two distinct sources (HoVer and Wikipedia summaries), both are derived from encyclopedic, formal domains, which may not fully represent the variety of sentence structures and informal language found in real-world use cases. Moreover, while all examples were systematically generated and reviewed, some bias may persist due to subjective decisions in the human review process. We attempt to mitigate this through cross-checking by an independent group of authors, but some residual bias may remain.
Thunder-NUBench primarily focuses on standard (sentence-level) negation and its distinction from local negation, contradiction, and paraphrase. Other important negation phenomena, such as double negation and negative polarity items (NPIs), are not directly addressed in this benchmark. Our current focus is on establishing a strong foundation for evaluating models’ understanding of standard negation. However, we aim to expand the evaluation to a broader range of negation phenomena in future work.
@@
This work does not involve the use of crowd-sourcing methods. Instead, all data included in the Thunder-NUBench benchmark has been carefully reviewed by the authors to ensure quality, relevance, and adherence to ethical standards. The datasets and tools used for training and evaluation are publicly available and used in compliance with their respective licenses.
@@
The Thunder-NUBench dataset is released under the CC BY-NC-SA 4.0 license, ensuring transparency, reproducibility, and accessibility for future research. We believe our work contributes positively to developing trustworthy and interpretable language models.
@@ Acknowledgments
We thank the anonymous reviewers and the meta-reviewer for their valuable feedback on this paper. We also sincerely thank Sungeun Hahm, Suyoung Park, Jongmin Kim, Yelim Ahn, Hyunji M. Park, Seorin Kim, and Jisoo Kim for their valuable contributions to the initial design of the negation dataset.
@@ Appendix A Typology of Contradiction
Table 7: Contradiction types from De Marneffe et al. (2008). Contradiction covers a broader scope than negation.
Contradictions in natural language can arise in diverse ways that go beyond simple negation. Following the typology of De Marneffe et al. (2008), contradictions can be grouped into seven categories: antonymy, explicit negation, numeric mismatch, factive/modal inconsistencies, structural reversals, lexical incompatibilities, and conflicts based on world knowledge. These categories reflect the fact that contradiction covers a broader semantic scope than negation alone. Table 7 summarizes these types with definitions and examples.
@@ Appendix B Copular Verbs
Copular verbs, also known as linking verbs, are verbs that connect the subject of a sentence to a subject complement, which can be a noun, adjective, or other expression that describes or identifies the subject. Unlike action verbs, copular verbs do not express actions but rather states or conditions. The most common copular verb in English is "to be" in its various forms (am, is, are, was, were). Other examples include "seem," "appear," "become," "feel," "look," "sound," "taste," and "smell" when used to describe the subject’s state (Hengeveld, 1986).
@@ Appendix C Negation of Implications
Negating implications presents challenges, as natural language intuitions often diverge from the rules of formal logic. Let’s say there is a conditional statement, "If I study hard, I will pass the bar exam." Formally, let $P$ denote "I study hard" and $Q$ denote "I will pass the bar exam." In classical logic, the conditional "if $P$, QQ" can be false only when $P$ is true and $Q$ is false. This implies that the negation of the conditional is $P\;\text{and}\;\text{Neg}(Q)$" ("I study hard and I won’t pass the bar exam,) while the conditional itself is equivalent to $\text{Neg}(P)\;\text{or}\;Q$" ("I don’t study hard or I will pass the exam) (Nguyen et al., 2023).
Psychological studies confirm that people often accept both "if $P$, Neg(QQ)" ("If I study hard, I won’t pass the exam.") and "if Neg($P$), QQ" ("If I don’t study hard, I will pass the exam."). However, the former can be interpreted as $\text{Neg}(P)\;\text{or}\;\text{Neg}(Q)$", and the latter $P\;\text{or}\;Q$", both of which are not equivalent to the original statement’s negation, $P\;\text{and}\;\text{Neg}(Q)$". "if Neg($P$), Neg(QQ)" ("If I don’t study hard, I won’t pass the exam.") is not the correct negation as well, as it is equivalent to $P\;\text{or}\;\text{Neg}(Q)$(Espino and Byrne, 2012).
@@ Appendix D Compound Sentences and Coordinating Conjunction
A compound sentence consists of two or more independent clauses joined by a coordinating conjunction. Each clause can stand alone, but they are combined to express related ideas (Gleitman, 1965).
@@ Appendix F Double Negation
Double negation refers to the use of two forms of grammatical negation within a single sentence. In standard English, only one negative form should be present in a subject-predicate construction; the presence of two negatives is generally considered non-standard and often results in an unintended meaning. For example, while "He’s going nowhere" is correct, "He’s not going nowhere" is ungrammatical. Another example is "I won’t bake no cake," which combines verb negation ("won’t") with object negation ("no cake"), resulting in a grammatically incorrect construction (Déprez et al., 2015).
In English, certain double negation constructions convey affirmative meanings rather than intensifying negation, effectively paraphrasing the original positive statement (e.g., $\neg\neg p\approx p$(Van der Wouden, 1996). This rhetorical device, known as litotes, often manifests in expressions such as "not bad," implying "good," or "not unhappy," implying "happy." Leveraging this phenomenon, we have generated paraphrase candidates for our dataset using such double negation patterns. For example,